How large a force is required to accelerate a 1400-kg car from rest to a speed of 30 m/s in a distance of 80 m?


1
Expert's answer
2021-03-21T11:37:28-0400

Let's first find the acceleration of the car:


"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{(30\ \dfrac{m}{s})^2}{2\cdot80\ m}=5.62\ \dfrac{m}{s^2}."

Finally, we can find the force from the Newton's Second Law of Motion:


"F=ma=1400\ kg\cdot5.62\ \dfrac{m}{s^2}=7868\ N."